OpenStudy (ttp):

Raven’s parents bought a 20’ ski boat for $22,000. After 2 years, the boat was worth 17,500. Using x as the number of years and y as the boat value, graph the two points. Part 1: [2 points] Describe the slope of the line which connects the two points. Part 2: [3 points] How can your graph be used to predict the value of the boat in 5 years? Part 3: [2 points] What is the projected boat value in 5 years?

OpenStudy (anonymous):

Point 1 = (0, 22000) Point 2 = (2, 17500) So far so good? 1. The slope is m = (y2 - y1) / (x2 - x1) Thus, m = (17500 - 22000) / (2 - 0) m = -2250 2. If you plot your points on a grid paper, you'll get a line.. and you can continue that line and predict the future value. That's why it's important for you to graph the points and do the exercise yourself. 3. What's the value for y on your line for x = 5? See your graph. It should be 10750..